Find a solution of p+q^2=0 passing through the line x0(s) =0, y0(s)=s, z0(s)=3s
1
Expert's answer
2020-12-25T13:57:57-0500

Let z=f(x+cy) be the trial solution of the given PDE

Let us take u= x+cy, so z=f(u)

Then we have p=dz/du, q=cdz/du

q/p=c

q=pc

p+p2c2=0 p(1+pc2)=0

1+pc2=0

p=-1/c2

dz=pdx+qdy=p(dx+cdy)=0

dx+cdy=0

dx=-cdy

x=-cy+c1

Substituting the value of p

p, we get

"dz=\frac{-(dx+cdy)}{c^2}" =0

z=-x/c2-y/c+c1


3s=-s/c+c1

s/c=c1-3s

c=s(c1-3s)

c1=3s+s/c

0=-cs+3s+s/c

0=-c2s+3cs+s

D=9s2 +4s

"c=\frac{-3s+\sqrt{9s^2+4s}}{-2}"


Answer: "\frac{4}{(-3s+\sqrt{9s^2+4s})^2}-(\frac{2}{(-3s+\sqrt{9s^2+4s})})^2=0"